# Characterization of flip process rules with the same trajectories

### v2

@@ 3.1. Key concepts and main lemmas
Our first lemma deals with twin vertices in the context of nearly induced copy functions between rooted graphs.
@@
The following hold for any nearly induced copy function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ from a rooted graph $\mathbf{F}$ to a rooted graph $\mathbf{H}$.
@@
Next, we describe a pair of complementary operations on rooted graphs. The first one shrinks clusters of twins in a rooted graph to obtain a twinfree version, while the second one blows up vertices of a (twinfree) rooted graph to clusters of twins. For a rooted graph $\mathbf{F}$ we define an equivalence relation $\sim_{\mathrm{tf}(\mathbf{F})}$ on $V(\mathbf{F})$ as follows. For $u,v\in V(\mathbf{F})$ we have $u\sim_{\mathrm{tf}(\mathbf{F})}v$ if $u=v$ or both $N(u)=N(v)$ and $|R(\mathbf{F})\cap\{u,v\}|\neq 1$ hold. Write $\mathcal{C}(\mathbf{F})$ for the set of equivalence classes of $\sim_{\mathrm{tf}(\mathbf{F})}$. For $v\in V(\mathbf{F})$ we write $v_{\mathrm{tf}(\mathbf{F})}$ for the unique equivalence class of $\sim_{\mathrm{tf}(\mathbf{F})}$ containing $v$. For $A\subseteq V(\mathbf{F})$ we set $A_{\mathrm{tf}(\mathbf{F})}=\{a_{\mathrm{tf}(\mathbf{F})}:a\in A\}$. The twinfree version of $\mathbf{F}$ is the rooted graph $\mathrm{tf}(\mathbf{F})$ on $\mathcal{C}(\mathbf{F})$ whose edges are the pairs $IJ$ for which there are $i\in I$ and $j\in J$ such that $ij\in E(\mathbf{F})$ and whose roots are the elements of $R(\mathbf{F})_{\mathrm{tf}(\mathbf{F})}$ with the linear order induced by $\leq_{R(\mathbf{F})}$. That $\mathrm{tf}(\mathbf{F})$ is twinfree follows easily from the definitions.
@@
For a rooted graph $\mathbf{F}$ and a vector $\mathbf{m}=(m_{i})_{i\in V(\mathbf{F})}\in\mathbb{N}_{0}^{V(\mathbf{F})}$ the $\mathbf{m}$-blowup of $\mathbf{F}$ is the rooted graph $\mathbf{F}(\mathbf{m})$ obtained as follows. First, replace each $i\in V(\mathbf{F})$ by an independent set $S_{i}$ of $m_{i}$ vertices and each edge $ij\in E(\mathbf{F})$ by a complete bipartite graph on $(S_{i},S_{j})$. We say that we blow up each vertex $i\in V(\mathbf{F})$ by a factor of $m_{i}$ to obtain $\mathbf{F}(\mathbf{m})$ from $\mathbf{F}$ and often omit mention of vertices $i\in V(\mathbf{F})$ where $m_{i}=1$. Next, let the set of roots be $\bigcup_{i\in R(\mathbf{F})}S_{i}$ ordered so that each $S_{i}$ forms an interval and for all $u\in S_{i}$ and $v\in S_{j}$ with $i\leq_{R(\mathbf{F})}j$ we have $u\leq_{R(\mathbf{F}(\mathbf{m}))}v$. Note that the linear order is not specified within each interval $S_{i}$; any order on each interval $S_{i}$ would give the same rooted graph up to isomorphism, so we conveniently avoid making these arbitrary choices. It follows easily from the definitions that every rooted graph $\mathbf{F}$ is isomorphic to the $\mathbf{m}$-blowup of its twinfree version $\mathrm{tf}(\mathbf{F})$ for some vector $\mathbf{m}$ of positive integers.
@@
Now we define notation and terminology to describe our quantification of induced densities at certain graphons. Let $(S,\leq_{S})$ be a finite ordered set with a partition $\mathcal{S}=\{S_{i}\}_{i\in I}$ indexed by a finite set $I$. Writing $a_{i}$ for the least element of $S_{i}$ according to $\leq_{S}$, the linear order $\leq_{I}$ on $I$ induced by $\leq_{S}$ is given by $i\leq_{I}j$ if and only if $a_{i}\leq_{S}a_{j}$. Denote by $Z(G)$ the collection of vectors $\mathbf{z}=(z_{i})_{i\in V(G)}\in[0,1]^{V(G)}$ satisfying $\sum_{i\in V(G)}z_{i}=1$. For $\mathbf{z}\in Z(G)$ write $\mathbf{\Omega}(\mathbf{z})$ for a partition $(\Omega_{i})_{i\in V(G)}$ of $\Omega$ such that $\Omega_{i}$ has measure $z_{i}$. For $\mathbf{z}\in Z(G)$ and $\mathbf{\Omega}(\mathbf{z})$ we say that the graphon $W_{G}^{\mathbf{z}}$ which is equal to $\mathbb{I}_{\left\{ij\in E(G)\right\}}$ on $\Omega_{i}\times\Omega_{j}$ is a $\mathbf{z}$-scaled graphon representation of $G$. For a graph $G$ and an ordered subset $X\subseteq V(G)$ the $X$-rooted version of $G$ is the rooted graph $\mathbf{F}$ obtained from $G$ by blowing up each vertex in $X$ by a factor of $2$ and designating as the ordered set of roots the set of duplicate vertices with the linear order induced by $\leq_{X}$. For $(x,y)\in\Omega^{2}$ the $(x,y)$-rooted version of $G$ for partition $(\Omega_{i})_{i\in V(G)}$ is the $X$-rooted version of $G$ where $X$ is the set $\{v\in V(G):\{x,y\}\cap\Omega_{v}\neq\varnothing\}$ ordered so that the first element $w$ satisfies $x\in\Omega_{w}$. We write $x_{\mathrm{rt}}$ (resp. $y_{\mathrm{rt}}$) for the root vertex whose twin $w$ in $X$ satisfies $x\in\Omega_{w}$ (resp. $y\in\Omega_{w}$).
@@
The following lemma gives an explicit quantification of induced densities at parametrized graphon representations of graphs. The availability of such an explicit quantification is central to our proof of Theorem 2.3.
@@
Suppose that $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}$ and fix $W\in\mathcal{W}$. By Theorem 2.2(i) we may pick trajectories $\Phi_{\mathcal{R}_{1}}^{\cdot}W$ and $\Phi_{\mathcal{R}_{2}}^{\cdot}W$ starting at $W$ which have maximal domains $\mathcal{D}_{\mathcal{R}_{1},W}$ and $\mathcal{D}_{\mathcal{R}_{2},W}$, respectively, that are open intervals containing 00. Let the function $f\colon\mathcal{D}_{\mathcal{R}_{1},W}\cup\mathcal{D}_{\mathcal{R}_{2},W}\to(\mathcal{W},\left\lVert\cdot\right\rVert_{\infty})$ be given by
(6) $f(t)=\begin{cases}\Phi_{\mathcal{R}_{1}}^{t}W&\textrm{if }t\in\mathcal{D}_{\mathcal{R}_{1},W}\\ \Phi_{\mathcal{R}_{2}}^{t}W&\textrm{if }t\in\mathcal{D}_{\mathcal{R}_{2},W}\setminus\mathcal{D}_{\mathcal{R}_{1},W}.\end{cases}$
Now the combination of (S3) and Definition 2.1 implies that $f$ is a trajectory starting at $W$ for both $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$, so by Theorem 2.2(i) we have $\mathcal{D}_{\mathcal{R}_{1},W}=\mathcal{D}_{\mathcal{R}_{2},W}$ and $\Phi_{\mathcal{R}_{1}}^{t}W=f(t)=\Phi_{\mathcal{R}_{2}}^{t}W$ for all $t\in\mathcal{D}_{\mathcal{R}_{1},W}$. Hence, $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ have the same trajectories. ∎
@@
Suppose that there is $W\in\mathcal{W}_{0}$ such that $\mathfrak{V}_{\mathcal{R}_{1}}W\neq\mathfrak{V}_{\mathcal{R}_{2}}W$. Setting $\varepsilon=\left\lVert\mathfrak{V}_{\mathcal{R}_{1}}W-\mathfrak{V}_{\mathcal{R}_{2}}W\right\rVert_{\infty}>0$, take $0<\eta\leq\frac{\varepsilon}{5C_{k}k(k-1)}$ and $U\in\mathcal{W}_{0}$ satisfying $\left\lVert U-W\right\rVert_{\infty}\in[0,\eta]$. By the triangle inequality for $\left\lVert\cdot\right\rVert_{\infty}$ and Lemmas 3.3 and 3.4, we have
$\varepsilon$ $=\left\lVert\mathfrak{V}_{\mathcal{R}_{1}}W-\mathfrak{V}_{\mathcal{R}_{2}}W\right\rVert_{\infty}$
$\leq\left\lVert\eta^{-1}(\Phi_{\mathcal{R}_{1}}^{\eta}U-\Phi_{\mathcal{R}_{2}}^{\eta}U)\right\rVert_{\infty}+\left\lVert\eta^{-1}(\Phi_{\mathcal{R}_{1}}^{\eta}U-U)-\mathfrak{V}_{\mathcal{R}_{1}}U\right\rVert_{\infty}+\left\lVert\eta^{-1}(\Phi_{\mathcal{R}_{2}}^{\eta}U-U)-\mathfrak{V}_{\mathcal{R}_{2}}U\right\rVert_{\infty}$
$\qquad+\left\lVert\mathfrak{V}_{\mathcal{R}_{1}}W-\mathfrak{V}_{\mathcal{R}_{1}}U\right\rVert_{\infty}+\left\lVert\mathfrak{V}_{\mathcal{R}_{2}}W-\mathfrak{V}_{\mathcal{R}_{2}}U\right\rVert_{\infty}$
$\leq\eta^{-1}\left\lVert\Phi_{\mathcal{R}_{1}}^{\eta}U-\Phi_{\mathcal{R}_{2}}^{\eta}U\right\rVert_{\infty}+2C_{k}k(k-1)\eta+2C_{k}\eta.$
Rearranging yields $\left\lVert\Phi_{\mathcal{R}_{1}}^{\eta}U-\Phi_{\mathcal{R}_{2}}^{\eta}U\right\rVert_{\infty}\geq\eta(\varepsilon-2C_{k}k(k-1)\eta-2C_{k}\eta)\geq\frac{\eta\varepsilon}{5}>0$, which implies $\Phi_{\mathcal{R}_{1}}^{\eta}U\neq\Phi_{\mathcal{R}_{2}}^{\eta}U$. Hence, $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ do not have the same trajectories. ∎
@@
Fix $q\in V^{\prime}(\mathbf{H})$ and set $S=V^{\prime}(\mathbf{H})$, $s=|S|$, $S^{\prime}=S\setminus\{q\}$ and $s^{\prime}=s-1$. Since $\mathbf{z}$ is subject to the condition $\sum_{i\in S}z_{i}=1$, we replace $z_{q}$ with $1-\sum_{i\in S^{\prime}}z_{i}$ in (8) to obtain an $s^{\prime}$-variable polynomial function $p$. The coefficient of each monomial $\prod_{i\in S^{\prime}}z_{i}^{m_{i}}$ in $p$ has the form $\sum_{\mathbf{n}}C_{\mathbf{m},\mathbf{n}}\cdot\alpha_{\mathbf{H}(\mathbf{n})}$ where the constants $C_{\mathbf{m},\mathbf{n}}$ satisfy $C_{\mathbf{m},\mathbf{m}}\neq 0$ and the sum is over all vectors $\mathbf{n}$ such that $n_{i}\leq m_{i}$ for all $i\in S^{\prime}$. Since $\alpha_{\mathbf{G}}\neq 0$, we may pick $\mathbf{p}$ such that $\alpha_{\mathbf{H}(\mathbf{p})}\neq 0$ and $\alpha_{\mathbf{H}(\mathbf{n})}=0$ for all $\mathbf{n}\neq\mathbf{p}$ where we have $n_{i}\leq p_{i}$ for all $i\in V(\mathbf{H})\setminus\{q\}$. The coefficient of $\prod_{i\in S^{\prime}}z_{i}^{p_{i}}$ in $p$ is nonzero, so $p$ is nonzero. Now $p$ is not identically zero on $[0,s^{-1}]^{s-1}$ by Lemma 3.5, so we may pick $\mathbf{y}\in Z(\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times})$ such that $\mathfrak{V}_{\mathcal{R}_{1}}W_{\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times}}^{\mathbf{y}}\neq\mathfrak{V}_{\mathcal{R}_{2}}W_{\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times}}^{\mathbf{y}}$. ∎

### v3

@@ 3.1. Key concepts and main lemmas
Our first lemma deals with twin vertices in the context of functions between rooted graphs that satisfy (F1) and (F2).
@@
The following hold for any function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ from a rooted graph $\mathbf{F}$ to a rooted graph $\mathbf{H}$ that satisfies (F1) and (F2).
@@
Next, we describe a pair of complementary operations on rooted graphs. The first one shrinks clusters of twins in a rooted graph to obtain a twinfree version, while the second one blows up vertices of a (twinfree) rooted graph to clusters of twins. Let $(S,\leq_{S})$ be a finite ordered set with a partition $\mathcal{S}=\{S_{i}\}_{i\in I}$ indexed by a finite set $I$. Writing $a_{i}$ for the least element of $S_{i}$ according to $\leq_{S}$, the linear order $\leq_{I}$ on $I$ induced by $\leq_{S}$ is given by $i\leq_{I}j$ if and only if $a_{i}\leq_{S}a_{j}$. For a rooted graph $\mathbf{F}$ we define an equivalence relation $\sim_{\mathrm{tf}(\mathbf{F})}$ on $V(\mathbf{F})$ as follows. For $u,v\in V(\mathbf{F})$ we have $u\sim_{\mathrm{tf}(\mathbf{F})}v$ if $u=v$ or both $N(u)=N(v)$ and $|R(\mathbf{F})\cap\{u,v\}|\neq 1$ hold. Write $\mathcal{C}(\mathbf{F})$ for the set of equivalence classes of $\sim_{\mathrm{tf}(\mathbf{F})}$. For $v\in V(\mathbf{F})$ we write $v_{\mathrm{tf}(\mathbf{F})}$ for the unique equivalence class of $\sim_{\mathrm{tf}(\mathbf{F})}$ containing $v$. For $A\subseteq V(\mathbf{F})$ we set $A_{\mathrm{tf}(\mathbf{F})}=\{a_{\mathrm{tf}(\mathbf{F})}:a\in A\}$. The twinfree version of $\mathbf{F}$ is the rooted graph $\mathrm{tf}(\mathbf{F})$ on $\mathcal{C}(\mathbf{F})$ whose edges are the pairs $IJ$ for which there are $i\in I$ and $j\in J$ such that $ij\in E(\mathbf{F})$ and whose roots are the elements of $R(\mathbf{F})_{\mathrm{tf}(\mathbf{F})}$ with the linear order induced by $\leq_{R(\mathbf{F})}$. That $\mathrm{tf}(\mathbf{F})$ is twinfree follows easily from the definitions.
@@
For a rooted graph $\mathbf{F}$ and a vector $\mathbf{m}=(m_{i})_{i\in V(\mathbf{F})}\in\mathbb{N}_{0}^{V(\mathbf{F})}$ the $\mathbf{m}$-blowup of $\mathbf{F}$ is the rooted graph $\mathbf{F}(\mathbf{m})$ obtained as follows. First, replace each $i\in V(\mathbf{F})$ by an independent set $S_{i}$ of $m_{i}$ vertices and each edge $ij\in E(\mathbf{F})$ by a complete bipartite graph on $(S_{i},S_{j})$. We say that we blow up each vertex $i\in V(\mathbf{F})$ by a factor of $m_{i}$ to obtain $\mathbf{F}(\mathbf{m})$ from $\mathbf{F}$ and often omit mention of vertices $i\in V(\mathbf{F})$ where $m_{i}=1$. Next, let the set of roots be $\bigcup_{i\in R(\mathbf{F})}S_{i}$ ordered so that each $S_{i}$ forms an interval and for all $u\in S_{i}$ and $v\in S_{j}$ with $i\leq_{R(\mathbf{F})}j$ we have $u\leq_{R(\mathbf{F}(\mathbf{m}))}v$. Note that the linear order is not specified within each interval $S_{i}$; any order on each interval $S_{i}$ would give the same rooted graph up to isomorphism, so we conveniently avoid making these arbitrary choices. It follows easily from the definitions that every rooted graph $\mathbf{F}$ with $r(\mathbf{F})\leq 2$ is isomorphic to the $\mathbf{m}$-blowup of its twinfree version $\mathrm{tf}(\mathbf{F})$ for some vector $\mathbf{m}$ of positive integers. Since in practice our rooted graphs have at most two roots, this is sufficient for our purposes; we remark that for many rooted graphs with at least three roots the statement above actually does not hold due to difficulties in handling the order on the roots.
@@
Now we define notation and terminology to describe our quantification of induced densities at certain graphons. Denote by $Z(G)$ the collection of vectors $\mathbf{z}=(z_{i})_{i\in V(G)}\in[0,1]^{V(G)}$ satisfying $\sum_{i\in V(G)}z_{i}=1$. For $\mathbf{z}\in Z(G)$ write $\mathbf{\Omega}(\mathbf{z})$ for a partition $(\Omega_{i})_{i\in V(G)}$ of $\Omega$ such that $\Omega_{i}$ has measure $z_{i}$. For $\mathbf{z}\in Z(G)$ and $\mathbf{\Omega}(\mathbf{z})$ we say that the graphon $W_{G}^{\mathbf{z}}$ which is equal to $\mathbb{I}_{\left\{ij\in E(G)\right\}}$ on $\Omega_{i}\times\Omega_{j}$ is a $\mathbf{z}$-scaled graphon representation of $G$. For a graph $G$ and an ordered subset $X\subseteq V(G)$ the $X$-rooted version of $G$ is the rooted graph $\mathbf{F}$ obtained from $G$ by blowing up each vertex in $X$ by a factor of $2$ and designating as the ordered set of roots the set of duplicate vertices with the linear order induced by $\leq_{X}$. For $(x,y)\in\Omega^{2}$ the $(x,y)$-rooted version of $G$ for partition $(\Omega_{i})_{i\in V(G)}$ is the $X$-rooted version of $G$ where the ordered subset $X=\{u,v\}\subseteq V(G)$ is such that $x\in\Omega_{u}$, $y\in\Omega_{v}$ and $u\leq_{X}v$. Note that we may have $u=v$; in this case, $X$ is a singleton set. Finally, we write $x_{\mathrm{rt}}$ (resp. $y_{\mathrm{rt}}$) for the root vertex whose twin $w$ in $X$ satisfies $x\in\Omega_{w}$ (resp. $y\in\Omega_{w}$).
@@
The following lemma gives an explicit quantification of induced densities at parametrized graphon representations of graphs. The availability of such an explicit quantification is central to our proof of Theorem 2.3. A function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ from a rooted graph $\mathbf{F}$ to a rooted graph $\mathbf{H}$ is induced copy if it satisfies (F1), (F2) and (F3).
@@
Suppose that $\mathfrak{V}_{\mathcal{R}_{1}}=\mathfrak{V}_{\mathcal{R}_{2}}$ on $\mathcal{W}$ and fix $W\in\mathcal{W}$. By Theorem 2.2(i) we have trajectories $\Phi_{\mathcal{R}_{1}}^{\cdot}W$ and $\Phi_{\mathcal{R}_{2}}^{\cdot}W$ starting at $W$ which have maximal domains $\mathcal{D}_{\mathcal{R}_{1},W}$ and $\mathcal{D}_{\mathcal{R}_{2},W}$, respectively, that are open intervals containing 0. The combination of (S3) and Definition 2.1 implies that $\Phi_{\mathcal{R}_{1}}^{\cdot}W$ is a trajectory starting at $W$ for $\mathcal{R}_{2}$ and $\Phi_{\mathcal{R}_{2}}^{\cdot}W$ is a trajectory starting at $W$ for $\mathcal{R}_{1}$, so by Theorem 2.2(i) $\Phi_{\mathcal{R}_{1}}^{\cdot}W$ and $\Phi_{\mathcal{R}_{2}}^{\cdot}W$ are restrictions of each other. Hence, we have $\mathcal{D}_{\mathcal{R}_{1},W}=\mathcal{D}_{\mathcal{R}_{2},W}$ and $\Phi_{\mathcal{R}_{1}}^{t}W=\Phi_{\mathcal{R}_{2}}^{t}W$ for all $t\in\mathcal{D}_{\mathcal{R}_{1},W}$, which means that $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ have the same trajectories. ∎
@@
Suppose that there is $W\in\mathcal{W}_{0}$ such that $\mathfrak{V}_{\mathcal{R}_{1}}W\neq\mathfrak{V}_{\mathcal{R}_{2}}W$. Setting $\varepsilon=\left\lVert\mathfrak{V}_{\mathcal{R}_{1}}W-\mathfrak{V}_{\mathcal{R}_{2}}W\right\rVert_{\infty}>0$, take $0<\eta\leq\frac{\varepsilon}{3C_{k}k(k-1)}$. By the triangle inequality for $\left\lVert\cdot\right\rVert_{\infty}$ and Lemmas 3.3 and 3.4, we have
$\varepsilon=\left\lVert\mathfrak{V}_{\mathcal{R}_{1}}W-\mathfrak{V}_{\mathcal{R}_{2}}W\right\rVert_{\infty}$ $\leq\left\lVert\eta^{-1}(\Phi_{\mathcal{R}_{1}}^{\eta}W-\Phi_{\mathcal{R}_{2}}^{\eta}W)\right\rVert_{\infty}+\sum_{i=1}^{2}\left\lVert\eta^{-1}(\Phi_{\mathcal{R}_{i}}^{\eta}W-W)-\mathfrak{V}_{\mathcal{R}_{i}}W\right\rVert_{\infty}$
$\leq\eta^{-1}\left\lVert\Phi_{\mathcal{R}_{1}}^{\eta}W-\Phi_{\mathcal{R}_{2}}^{\eta}W\right\rVert_{\infty}+2C_{k}k(k-1)\eta.$
Rearranging yields $\left\lVert\Phi_{\mathcal{R}_{1}}^{\eta}W-\Phi_{\mathcal{R}_{2}}^{\eta}W\right\rVert_{\infty}\geq\eta(\varepsilon-2C_{k}k(k-1)\eta)\geq\frac{\eta\varepsilon}{3}>0$, which implies $\Phi_{\mathcal{R}_{1}}^{\eta}W\neq\Phi_{\mathcal{R}_{2}}^{\eta}W$. Hence, $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ do not have the same trajectories. ∎
@@
Fix $q\in V^{\prime}(\mathbf{H})$ and set $S=V^{\prime}(\mathbf{H})$, $s=|S|$, $S^{\prime}=S\setminus\{q\}$ and $s^{\prime}=s-1$. Since $\mathbf{z}$ is subject to the condition $\sum_{i\in S}z_{i}=1$, we replace $z_{q}$ with $1-\sum_{i\in S^{\prime}}z_{i}$ in (7) to obtain an $s^{\prime}$-variable polynomial function $p$. The coefficient of each monomial $\prod_{i\in S^{\prime}}z_{i}^{m_{i}}$ in $p$ has the form $\sum_{\mathbf{n}}C_{\mathbf{m},\mathbf{n}}\cdot\alpha_{\mathbf{H}(\mathbf{n})}$ where the constants $C_{\mathbf{m},\mathbf{n}}$ satisfy $C_{\mathbf{m},\mathbf{m}}\neq 0$ and the sum is over all vectors $\mathbf{n}$ such that $n_{i}\leq m_{i}$ for all $i\in S^{\prime}$. Since $\alpha_{\mathbf{G}}\neq 0$, we may pick $\mathbf{p}$ such that $\alpha_{\mathbf{H}(\mathbf{p})}\neq 0$ and $\alpha_{\mathbf{H}(\mathbf{n})}=0$ for all $\mathbf{n}\neq\mathbf{p}$ where we have $n_{i}\leq p_{i}$ for all $i\in V(\mathbf{H})\setminus\{q\}$. For example, we may start with $\mathbf{p}$ such that $\mathbf{G}=\mathbf{H}(\mathbf{p})$, replace the current $\mathbf{p}$ with any $\mathbf{n}\neq\mathbf{p}$ such that $\alpha_{\mathbf{H}(\mathbf{n})}\neq 0$ and $n_{i}\leq p_{i}$ for all $i\in V(\mathbf{H})\setminus\{q\}$ if possible, and take the vector $\mathbf{p}$ at termination. The coefficient of $\prod_{i\in S^{\prime}}z_{i}^{p_{i}}$ in $p$ is the nonzero $C_{\mathbf{p},\mathbf{p}}\cdot\alpha_{\mathbf{H}(\mathbf{p})}$, so $p$ is nonzero. Now $p$ is not identically zero on $0,s^{-1}]^{s-1}$ by Lemma 3.5, so we may pick $\mathbf{y}\in Z(\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times})$ such that $\mathfrak{V}_{\mathcal{R}_{1}}W_{\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times}}^{\mathbf{y}}\neq\mathfrak{V}_{\mathcal{R}_{2}}W_{\mathrm{tf}(\mathbf{G}^{\varnothing})^{\times}}^{\mathbf{y}}$. ∎